OpenStudy (anonymous):

30v^4w^3 - 12v^4w^3 = ??? HELP?

OpenStudy (anonymous):

\[(30v^{4}w ^{3})-(12v ^{4}w ^{3})\]

OpenStudy (anonymous):

Is that you equation?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

17v^4w^3 should be the answer i believe

OpenStudy (anonymous):

No...that not it

OpenStudy (anonymous):

I need help with how to do it, not the answer

OpenStudy (anonymous):

yeah ok pretty easy ill explain it

OpenStudy (anonymous):

Kay, i just got confused.

OpenStudy (anonymous):

I can do this one :)

OpenStudy (anonymous):

You have: \[(30v^4w^3)-(12v^4w^3)\] Pick a letter, let's say \(a\) for amber haha Let \(a=v^4w^3\) So then you can simplify: \[(30v^4w^3)-(12v^4w^3)=30a-12a=18a\] And we can substitute back: \[18a=18v^4w^3\]

OpenStudy (anonymous):

\[(30v ^{4}w ^{3}-12v ^{4}-w ^{3})\]

OpenStudy (anonymous):

See how its a negative 12v^4 and negative w^3 bceause you multyply the -(12v^4W^3)

OpenStudy (anonymous):

Reaper, \[-12v^4w^3\neq-12v^4-w^3\]...

OpenStudy (anonymous):

ohh shoot my bad yes your right i was thining smth else

OpenStudy (anonymous):

I was told to just line them up vertically and subtract by my teacher

OpenStudy (anonymous):

If you had \(-(12v^4+w^3)\), then you'd be correct :)

OpenStudy (anonymous):

yeah keith is right

OpenStudy (anonymous):

that works to amber

OpenStudy (anonymous):

Ohh you could do that too! Like so: